### v1

# Large-Scale Calculations of $\beta$-Decay Rates and Implications for $r$-Process Nucleosynthesis

Abstract
Nuclear $\beta$ decay is a key element of the astrophysical rapid neutron capture process ($r$-process). In this paper, we present state-of-the-art global $\beta$-decay calculations based on the quantified relativistic nuclear energy density functional theory and the deformed proton-neutron quasiparticle random-phase approximation. Our analysis considers contributions from allowed and first-forbidden transitions. We used two point-coupling functionals with carefully calibrated time-odd terms and isoscalar pairing strength. The new calculations display consistent results for both employed functionals, especially near the neutron drip line, suggesting slower $\beta$ decays past the $N=126$ neutron shell closure than in commonly used $\beta$-decay models. The new rates, along with the existing rates based on the latest non-relativistic calculations, are found to slow down the synthesis of heavy elements in the $r$-process and significantly reduce the contribution of neutron-induced fission.
1 Introduction
The rapid neutron capture process ($r$-process) is responsible for synthesizing the heaviest elements that naturally exist in the Universe (Burbidge et al., 1957; Cameron, 1957; Horowitz et al., 2019; Cowan et al., 2021; Arcones & Thielemann, 2022). Although its exact site remains elusive, it is expected to occur in explosive and neutron-rich astrophysical events, such as compact binary mergers and rare types of core-collapse supernovae. The process involves thousands of species of neutron-rich nuclei, many of which have yet to be experimentally studied. They undergo various nuclear reactions and decays, including neutron capture, photodissociation, $\beta$ decay, and fission. As most of the nuclei involved are beyond the current reach of experiments, theoretical description of their properties has a profound impact on our understanding of the $r$-process. As a step towards consistent evaluation of the required nuclear physics input with controlled extrapolation, this work focuses on the influence of $\beta$ decay. The basic role of $\beta$ decay in the $r$-process is to transfer the material from the isotopic chain with proton number $Z$ to the one with $Z+1$. During the $(n,\gamma)\leftrightarrow(\gamma,n)$ equilibrium, the most abundant isotope in a given isotopic chain, so-called the “waiting-point” nucleus, effectively determines the $\beta$-decay rate of the isotopic chain. The half-lives of waiting-point nuclei with neutron magic numbers $N=82$ and $126$ qualitatively explain the origin of the second and third $r$-process abundance peaks at mass numbers $A\sim 130$ and $196$, respectively. After the $(n,\gamma)\leftrightarrow(\gamma,n)$ equilibrium breaks down, $\beta$ decay, as well as $\beta$-delayed neutron emission, neutron capture, and fission, are the main mechanisms that affect the final abundance pattern (Surman et al., 1997; Arcones & Martínez-Pinedo, 2011; Mumpower et al., 2012a, b, c; Goriely et al., 2013; Eichler et al., 2015). Fission determines the end point of the $r$-process path, past $Z\sim 90$ (Panov & Thielemann, 2003; Martínez-Pinedo et al., 2007; Panov, I. V. et al., 2010; Eichler et al., 2015; Giuliani et al., 2018; Mumpower et al., 2018; Vassh et al., 2019; Giuliani et al., 2020), redistributing and increasing the relative abundance of heavy nuclei. This mechanism is referred to as “fission (re)cycling”. Neutron-induced fission has been discussed as the dominant fission channel (Panov et al., 2005; Martínez-Pinedo et al., 2007; Petermann et al., 2012); however, depending on the time scale and amount of the population of fissile nuclei, the contribution of neutron-induced fission may be significantly altered.
In the majority of $r$-process studies, $\beta$-decay data have been taken from the phenomenological Finite Range Droplet Model (FRDM), combined with the quasiparticle random phase approximation (QRPA) (Möller et al., 1997, 2003, 2019). More microscopic approaches, whether based on the nuclear shell model (Sharma et al., 2022; De Gregorio et al., 2024) or ab initio methods (Gysbers et al., 2019), remain limited due to the unfavorable scaling of problem complexity with system size. Leveraging the balance between scalability and rigorous microscopic description, the energy density functional (EDF) theory is a suitable many-body method for providing reliable predictions of nuclear properties for $r$-process studies, as it is computationally tractable across the entire chart of nuclides. Starting from the nuclear ground state, excitations can be obtained by considering the linear response of the nucleus within QRPA. Until recently, most large-scale calculations of $\beta$-decay rates based on the nuclear EDF framework and QRPA have been restricted to spherical nuclei (Engel et al., 1999; Niu et al., 2013; Marketin et al., 2016). Only a single global calculation has considered axially-deformed nuclei (Ney et al., 2020). Since global sets of theoretical $\beta$-decay rates are scarce, the literature on the sensitivity of $r$-process simulations on $\beta$ decay has been limited (Niu et al., 2013; Mumpower et al., 2014; Lund et al., 2023; Kullmann et al., 2023; Hao et al., 2023), with the primary focus on the $r$-process observables such as the abundance pattern and electromagnetic emission. The dynamic effect of $\beta$ decay on the evolution of other reactions, decays, and abundances has rarely been explicitly discussed. In particular, the impact of $\beta$-decay modeling on the fission contribution has been largely unexplored.
In this Letter, we present new global $\beta$-decay calculations using the recently developed axially-deformed relativistic QRPA framework Ravlić et al. (2024), with the residual interaction optimized in Ravlić & Nazarewicz (2025). Using this tool, we investigate the effect of the $\beta$ decay on the $r$-process dynamics, through comparisons with other global models.
2 Method
2.1 Calculation of $\beta$-decay half-lives
The ground-state of the nucleus is obtained with the relativistic Hartree-Bogoliubov theory assuming axially-deformed and reflection-invariant nuclear shapes (Nikšić et al., 2014). Two point-coupling interactions are employed in this work: DD-PC1 (Nikšić et al., 2008) and DD-PCX (Yüksel et al., 2019), both adopting separable pairing interaction (Tian et al., 2009). The ground-state quadrupole deformation is obtained by performing constrained calculations with the quadrupole deformation $\beta_{2}\in[-0.5,0.5]$, and subsequently releasing the constraint to find the configuration that minimizes the total energy. Odd-$A$ and odd-odd nuclei are treated by self-consistent quasiparticle blocking within the equal-filling approximation (Perez-Martin & Robledo, 2008). Calculations are performed by discretizing quasiparticle states in a stretched axially-deformed harmonic oscillator basis. We use $N_{\rm osc}=18$ shells for nuclei up to $Z=103$ and for heavier nuclei $N_{\rm osc}=20$. To obtain excited states in the daughter nucleus, we employ the linear response approach, developed in Ravlić et al. (2024). The residual interaction couplings, including the isoscalar pairing interaction strength $V_{0}^{is}$, and Landau-Migdal coupling $g_{0}$, in addition to effective axial-vector coupling $g_{A}$, are used as optimized in the recent work (Ravlić & Nazarewicz, 2025). The $\beta$-decay rates $\lambda_{\beta}$ are calculated by the contour integration around a suitably chosen contour $\mathcal{C}$ encircling the $Q_{\beta}$ window:
$\lambda_{\beta}=\frac{\text{ln}2}{K}\frac{1}{2\pi i}\oint_{\mathcal{C}}d\omega\int\limits_{1}^{W_{0}[\omega]}dWf(W,Z)C(W,\omega),$ (1)
where the phase-space factor is
$f(W,Z)=pW(W_{0}[\omega]-W)^{2}F(Z,W),$ (2)
with $W$ being the electron energy (in $m_{e}c^{2}$ unit), $p=\sqrt{W^{2}-1}$ – its momentum (in $m_{e}c$), and $F(Z,W)$ being the Fermi function, taking into account the distortion of the electron wave function. The end-point energy is given by
$W_{0}[\omega](m_{e}c^{2})=\Delta_{np}-\omega+\lambda_{n}-\lambda_{p},$ (3)
where $\lambda_{n(p)}$ is the neutron(proton) chemical potential, and $\Delta_{np}=1.293$ MeV is the neutron-proton mass difference. The nuclear matrix elements are contained within the shape-factor
$C(W,\omega)=k(\omega)+ka(\omega)W+\frac{kb(\omega)}{W}+kc(\omega)W^{2},$ (4)
where $k,ka,kb,$ and $kc$ are tabulated in Behrens & Bühring (1971); Bambynek et al. (1977). The super-allowed decay constant is taken as $K=6137$ s. The circular contour is discretized with a 30-point Gauss-Laguerre quadrature. The integration over electron energy $W$ is performed by Lagrange interpolation of the integrand on a 20-point Chebyshev grid, as discussed in Ravlić & Nazarewicz (2025). Subsequently, the half-lives are obtained as $T_{1/2}=\text{ln}2/\lambda_{\beta}$. We include the allowed Gamow-Teller (GT) $1^{+}$ transitions, as well as the first-forbidden (FF) $0^{-},1^{-}$, and $2^{-}$ $\beta$ decays. Calculations are performed in the range of nuclei between $Z=20$ and $Z=110$, containing 4598 individual nuclides. For each isotopic chain, calculations start close to the valley of stability, and end at the two-neutron drip-line theoretically estimated in Ravlić et al. (2023).
Table 1: Summary of global $\beta$-decay models considered in this work for $r$-process simulations, including the method used to compute the rates, restrictions on nuclear shapes, and contribution of different multipoles. GT corresponds to $1^{+}$ and $F$ to $0^{-},1^{-}$ and $2^{-}$ multipoles.
Model Method Geometry Multipoles
FRDM Möller et al. (2019) micro-macro + QRPA(schematic) deformed GT+FF(gross)
D3C Marketin et al. (2016) EDF(covariant) + QRPA(matrix) spherical GT+$F$
SkO Ney et al. (2020) EDF(Skyrme) + QRPA(linear response) deformed GT+$F$
DD-PC1 and DD-PCX (this work) EDF(covariant) + QRPA(linear response) deformed GT+$F$
2.2 Nuclear reaction network
The $r$-process abundance evolution is calculated with the nuclear reaction network code PRISM (Sprouse, 2020). In addition to our new DD-PC1 and DD-PCX sets of $\beta$-decay rates, a summary of $\beta$-decay models employed in this work is shown in Table 1. In particular, it includes the FRDM + QRPA $\beta$-decay tables from Möller et al. (2003, 2019), based on a schematic separable QRPA interaction and phenomenological estimates of FF transitions. Other rates based on the EDF framework include the predictions by Ney et al. (2020), calculated with the axially-deformed charge-exchange linear response QRPA with the Skyrme SkO’ functional, and relativistic matrix-QRPA results with the EDF D3C* from Marketin et al. (2016), restricted to spherical nuclei. Other nuclear properties, such as one-neutron separation energies and fission data, are primarily based on the FRDM model and experimental data. Separate calculations were also performed with the REACLIB nuclear data library (Cyburt et al., 2010). See the Supplementary Material for the details and results (Ravlić et al., 2025).
For astrophysical conditions, we first employ the parameterized wind model of Panov & Janka (2009), with an initial entropy per baryon of $s/k_{B}=40$, and an expansion timescale of 20 ms. The initial electron fraction $Y_{e}$, a measure of the neutron-richness of the environment, ranges from 0.15 to 0.35 in 0.025 intervals. The initial temperature of these astrophysical trajectories is 10 GK, and the initial compositions of the nuclei were calculated using the SFHo equation of state (Steiner et al., 2013). Additionally, we also use a trajectory with an initial condition with $s/k_{B}=20$ and $Y_{e}=0.05$, to simulate a neutron-rich environment in which fission cycling is expected to occur.
To investigate the effect of the new $\beta$-decay rates in more realistic astrophysical conditions, we consider thermodynamical evolution from two sets of astrophysical simulations: a magneto-rotational supernova (Reichert et al., 2021a; Obergaulinger & Aloy, 2017; Reichert et al., 2021b), in which it has been suggested that the production of nuclei past the third $r$-process peak at $A\sim 195$ occurs; and outflow from the accretion disk formed after a merger of binary neutron stars (Metzger & Fernández, 2014; Holmbeck et al., 2019). Similarly to the case of the parametrized trajectories, nucleosynthesis calculations were initiated at $T=10$ GK, and the composition at the time was calculated using the SFHo equation of state.
3 Results and Discussion
Figure 1: Comparison between (a) mean averaged half-lives $\bar{r}$ and (b) their spread $\sigma$ obtained in this work ($D$-PC1 and $D$-PCX) with the results of SkO’ model (Ney et al., 2020), D3C* (Marketin et al., 2016), and FRDM (Möller et al., 2003). The experimental data are taken from  (Kondev et al., 2021). Averaging in Eq. (5) is performed over the nuclei with half-lives up to the maximum experimental half-life $T_{1/2}^{\rm exp,max}$ in the range from $10^{-1}$ s to $10^{4}$ s.
Figure 2: Comparison of theoretical $\beta$-decay rates relative to predictions of our $D$-PC1 EDF model: (a) our $D$-PCX model, (b) SkO’ model (Ney et al., 2020), (c) D3C* model (Marketin et al., 2016), and (d) FRDM model (Möller et al., 2019).
3.1 Benchmarking theoretical rate calculations
As a first step in validating new rate calculations, it is important to compare the results with available data from NUBASE2020 (Kondev et al., 2021). Since the half-lives span orders of magnitude, it is common to employ the following quantifiers for the mean value:
$r_{i}=\log_{10}\frac{T_{1/2}^{\text{calc.}(i)}}{T_{1/2}^{\text{exp.}(i)}},\quad\bar{r}=\frac{1}{N}\sum\limits_{i=1}^{N}r_{i},$ (5)
for a sample of $N$ nuclei with measured $T_{1/2}^{\rm exp.}$ and calculated $T_{1/2}^{\rm calc.}$ half-lives. The spread around the mean value is defined as
$\sigma=\sqrt{\frac{1}{N}\sum\limits_{i=1}^{N}(r_{i}-\bar{r})^{2}}.$ (6)
Results of the inter-model comparison are shown in Fig. 1, by restricting the summations in Eqs. (5) and (6) to nuclei with experimental half-lives below $T_{1/2}^{\rm exp,max}$, i.e., $T_{1/2}^{\text{exp.}(i)}<T_{1/2}^{\rm exp,max}$. Our DD-PC1 and DD-PCX results are compared to SkO’ (Ney et al., 2020), D3C* Marketin et al. (2016), and FRDM (Möller et al., 2003, 2019) predictions. In this comparison, we only consider nuclei with $T_{1/2}^{\rm calc.}<10^{6}$ s. As shown in Fig. 1(a), all models tend to overestimate the experimental half-lives. Both DD-PC1 and DD-PCX show a similar pattern, with the maximum discrepancy around $T_{1/2}^{\rm exp,max}=10$ s. In particular, DD-PCX shows excellent accuracy at $T_{1/2}^{\rm exp,max}=10^{4}$ s, with the mean relative error under 5%. For nuclei with shorter experimental half-lives, the maximum discrepancy is 20% for DD-PCX and 32% for DD-PC1. Overall, the mean error of our models is similar to other global theoretical models. In Fig. 1(b), we show the spread of the predictions around the mean, Eq. (6). Here, all models display a universal trend where $\sigma$ gradually increases with $T_{1/2}^{\rm exp,max}$.
In Fig. 2 we compare different theoretical $\beta$-decay rates, across the nuclear landscape relative to DD-PC1. Overall, both EDFs from this work show excellent agreement with SkO’. Discrepancies are seen closer to the valley of stability, where half-lives are quite long, meaning that the rates are determined by only a few peaks of the strength function within the $Q_{\beta}$-window. Otherwise, regions of somewhat larger discrepancies exist around magic nuclei with $N=50,Z=28$, $N=82,Z=50$, and $N=184$. While our rates tend to be smaller than SkO’ for these nuclei, the differences are within an order of magnitude. The excellent agreement with SkO’ towards the neutron drip-line is particularly remarkable, considering significant model extrapolations involved. Additionally, both EDFs used in this work show very consistent predictions, indicating a relatively small systematic uncertainty due to the choice of a functional. On the other hand, the D3C* spherical model exhibits significant discrepancies for $Z>60$ and $N>126$. We note that it is difficult to assess whether the differences solely originate from the assumption of sphericity in D3C* or due to the specifics of effective interaction. Somewhat smaller differences are obtained for FRDM, around one order of magnitude, being especially pronounced around and above the $Z=82$ shell closure. The half-lives predicted in this work, $Q_{\beta}$ values, and other ground-state properties are available in a public repository (Saito et al., 2025).
Figure 3: Final abundance patterns (top), the evolution of abundance-averaged proton number $\langle Z\rangle_{Y}$ as a function of time (middle), and the fission flow and neutron abundance as functions of time (bottom), for the s20ye0050 (left column) and s40ye0175 (right column) trajectories, respectively. The scaled solar abundance pattern (Goriely, 1999) is shown for reference.
3.2 Impact on the $r$-process
We have performed nuclear reaction network calculations employing the new DD-PC1 and DD-PCX values of $\beta$-decay rates and the existing predictions of SkO’, FRDM, and D3C*. Overall, we find that the $r$-process calculations using the DD-PC1 and DD-PCX $\beta$-decay rates are generally consistent with the SkO’ calculations, and are significantly different from those with FRDM and D3C*.
Let us define the flow $F_{P}$ of a process (that is, decay or reaction) $P$:
$F_{P}(t)\equiv\sum_{\text{all nuclei}}\lambda_{P}(t)\rho(t)^{N-1}\prod_{i=R_{1},\ldots,R_{N}}Y(Z_{i},A_{i},t),$ (7)
where $\lambda_{P}(t)$ is the rate of the process $P$ at time $t$, $\rho(t)$ is the baryon density of the environment at time $t$, $R_{1},\ldots,R_{N}$ are the reactants of the process $P$, and $Y(Z,A,t)$ is the abundance of the reactant $(Z,A)$.
Since the results obtained with the nuclear properties taken from the REACLIB nuclear data library and from FRDM are very similar, in the following we will only show the results based on FRDM data in the main text. The complete set of results can be found in the Supplementary Material (Ravlić et al., 2025).
3.2.1 Parameterized trajectories
We use ten parameterized trajectories to systematically assess the effect of different $\beta$-decay predictions. Our calculations show that $\beta$-decay rates significantly impact the abundances above the second $r$-process peak at $A\sim 130$. Therefore, to highlight this effect, we focus on trajectories corresponding to neutron-rich conditions s20ye0050 ($s/k_{B}=20$ and $Y_{e}=0.05$) and s40ye0175 ($s/k_{B}=40$ and $Y_{e}=0.175$), where $s/k_{B}$ and $Y_{e}$ are the initial entropy per baryon and electron fraction, respectively. s20ye0050 approximates an extremely neutron-rich condition, such as dynamical ejecta from a neutron star merger, where fission cycling is expected to occur, whereas s40ye0175 is a moderately neutron-rich condition. The complete set of results for ten trajectories can be found in Ravlić et al. (2025).
In Fig. 3(a,b), most notable effects of $\beta$ decays on the final abundance patterns can be seen around $140\leq A\leq 170$ for s20ye0050 and $130\leq A\leq 180$ for s40ye0175, respectively. Fig. 3(c,d) shows that the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates consistently lead to a slower increase in the average proton number, compared to FRDM or D3C*. Since the neutron density of the environment decreases as the ejecta expand over time, the delayed population of fissile nuclei also delays the fission cycling process, primarily driven by neutron-induced fission and subsequent neutron capture and $\beta$ decay. In an extremely neutron-rich condition s20ye0050, a significant number of neutron-induced fission events still occur, as shown in Fig. 3(e). However, compared to D3C* and FRDM, subsequent neutron capture on the fission fragments would be hindered by the delayed timing of neutron-induced fission, leading to a dip in abundance patterns around $140\leq A\leq 170$ with DD-PC1, DD-PCX, and SkO’. In the case of s40ye0175, Fig. 3(f) shows that the delayed population of fissile nuclei leads to a significant decrease in the flow of neutron-induced fission. This results in a significant decrease in abundance below $A\sim 180$ due to the absence of fission cycling. While neutron-induced fission has been considered to be the dominant channel in the $r$-process nucleosynthesis (Panov et al., 2005; Martínez-Pinedo et al., 2007; Petermann et al., 2012), our calculations with the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates suggest that the flow of neutron-induced fission is comparable to, or only slightly larger than, that of $\beta$-delayed fission (see Ravlić et al. (2025)). Although for the fission yields we assume a 50/50 split, as the focus of the work is the new $\beta$-decay rates, in future studies, for a more detailed assessment of the effect of fission yields, we plan to obtain realistic fission yields using the consistent EDF approach (Flynn et al., 2022; Lay et al., 2024).
3.2.2 Trajectories from Astrophysical Simulations
We consider two astrophysical simulations to provide more realistic $r$-process thermodynamic conditions: a mangeto-rotational supernova (Obergaulinger & Aloy, 2017; Reichert et al., 2021a), which we refer to as “MRSN”, and accretion disk outflow from a binary neutron star merger (Metzger & Fernández, 2014; Holmbeck et al., 2019), which we refer to as “disk outflow”. Each case consists of a set of ten thermodynamic trajectories, and the results in select trajectories will be highlighted. The complete results of the nuclear reaction network calculations are given in  Ravlić et al. (2025).
MRSN has been studied as a site that can produce elements up to the third $r$-process peak in jet-like ejecta along the axis of the rotation under strong magnetic fields, by escaping the exposure to neutrinos, which would make the ejecta less neutron-rich (Mösta et al., 2014, 2015, 2018; Halevi & Mösta, 2018; Cameron, 2003; Reichert et al., 2022; Winteler et al., 2012; Nishimura et al., 2006, 2015, 2017; Symbalisty et al., 1985; Reichert et al., 2021a).
Figure 4(a) shows the final abundance patterns for the most neutron-rich condition among the MRSN trajectories, and this is the only trajectory capable of synthesizing elements up to the third $r$-process peak. With the DD-PC1, DD-PCX, and SkO’ $\beta$-decay rates, the third $r$-process peak is reduced, compared to FRDM or D3C*. This indicates that $\beta$-decay rates, especially in the vicinity of and past the $N=126$ neutron shell closure, compete with the expansion timescale of the ejecta and substantially affect the nucleosynthetic outcome. The calculations with the REACLIB nuclear data library in Ravlić et al. (2025) yield more pronounced third $r$-process peaks, however, DD-PC1, DD-PCX, and SkO’ still consistently predict smaller peaks than FRDM and D3C*. These results suggest that conclusions on the production of elements in the third $r$-process peak and above in the MRSN conditions may be substantially dependent on nuclear physics inputs for the nucleosynthesis calculations, in particular $\beta$-decay rates.
Figure 4: Final abundance patterns with different sets of $\beta$-decay rates, for (a) the most neutron-rich condition among the MRSN trajectories, and (b) the most neutron-rich condition among the disk outflow trajectories. The panel (c) shows the temporal evolution of the flow of neutron-induced fission and neutron abundance $Y_{n}$ for the disk outflow condition (b).
For the disk outflow condition, we highlight the most neutron-rich trajectory with the initial entropy $s/k_{B}=21$ and electron fraction $Y_{e}=0.123$. The final abundances and the temporal evolution of the flow of neutron-induced fission and neutron abundance, calculated with different sets of $\beta$-decay tables combined with the FRDM-based nuclear properties, are shown in Fig. 4(b,c). Similarly to the case of the parameterized trajectory s40ye0175, there is an orders-of-magnitude decrease in neutron-induced fission flows in DD-PC1, DD-PCX, and SkO’, compared to D3C* and FRDM. With the former models, we find that the flows of $\beta$-delayed fission and neutron-induced fission are comparable. The reduced neutron-induced fission flow leads to an underproduction of the second $r$-process peak around $A\sim 130$ and accumulation of abundances around $200\leq A\leq 210$.
Holmbeck et al. (2019), who considered similar astrophysical conditions, have suggested that most actinide enrichments in metal-poor $r$-process enhanced stars are explained by an $r$-process source with a non-dominant amount of very neutron-rich, fission-cycling component. Given the significant effect of $\beta$ decay on fission flows indicated by our results, the understanding of actinide enrichment, in particular the distribution of the neutron-richness ($Y_{e}$) of the ejecta, is also likely affected by the choice of $\beta$-decay rates. Finally, the effect of $\beta$ decay on spontaneous fission flows is also present for the parameterized and disk outflow trajectories, consistent with the findings in Lund et al. (2023).
4 Conclusions
In this Letter, we presented the state-of-the-art global predictions of $\beta$-decay rates based on the relativistic nuclear EDF theory with DD-PC1 and DD-PCX functionals, including the effect of axial deformation of nuclei. The new rates agree well with the deformed non-relativistic SkO’ EDF predictions (Ney et al., 2020). These $\beta$-decay rates suggest slower $\beta$ decays, especially past the $N=126$ neutron shell closure, and are distinctly different from the phenomenological FRDM (Möller et al., 2019) or spherical relativistic EDF D3C* (Marketin et al., 2016) tabulations, which have been commonly used in $r$-process simulations. This significantly reduces the flow of neutron-induced fission, which affects the final abundance patterns. In less neutron-rich environments, such as magneto-rotational supernovae, the $\beta$-decay rates may compete with the expansion timescale of the ejecta, thereby affecting heavy nucleosynthesis. Since most astrophysical analyses depend on specific global $\beta$-decay models, such a choice may affect our understanding of the astrophysical environment of the $r$-process. The results presented in this work constitute a step in our effort towards comprehensive theoretical predictions based on the quantified nuclear EDF framework, which will enable a controlled assessment of uncertainties due to different nuclear properties.
5 Acknowledgments
This work was supported by the U.S. Department of Energy under Award No. DOE-DE-NA0004074 (NNSA, the Stewardship Science Academic Alliances program), the DOE Office of Science under Grants DE-SC0013365 and DE-SC0023175 (Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing), and by the U.S. National Science Foundation under grant number 21-16686 (NP3M). Computational resources were provided in part by the Institute for Cyber-Enabled Research at Michigan State University. This work used TAMU ACES at TAMU through allocation PHY250120 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants No. 2138259, 2138286, 2138307, 2137603, and 2138296.